### v1

# Thermal transport through a single trapped ion under strong laser illumination

Abstract
In this work, we study quantum heat transport in a single trapped ion, driven by laser excitation and coupled to thermal reservoirs operating at different temperatures. Our focus lies in understanding how different laser coupling scenarios impact the system dynamics. As the laser intensity reaches a regime where the ion’s electronic and motional degrees of freedom strongly couple, traditional approaches using phenomenological models for thermal reservoirs become inadequate. Therefore, the adoption of the dressed master equation (DME) formalism becomes crucial, enabling a deeper understanding of how distinct laser intensities influence heat transport. Analyzing the heat current within the parameter space defined by detuning and coupling strength, we observe intriguing circular patterns which are influenced by the ion’s vibrational frequency and laser parameters, and reveal nuanced relationships between heat transport, residual coherence, and system characteristics. Our study also reveals phenomena such as negative differential heat conductivity and asymmetry in heat current flow, offering insights into the thermal properties of this essential quantum technology setup.
I Introduction
The desire to build universal quantum computers nielsen2010 has led to a remarkable development of methods for controlling quantum systems and, additionally, the creation of increasingly smaller and more complex quantum devices. In order to have the level of control necessary for this application, it is important, among other things, to understand how these quantum systems interact with their environment breuer2002 . One of the efforts in this direction is that of understanding nonequilibrium processes, e.g., transport of energy, and how the response changes according to system or reservoir properties.
On the other hand, few quantum technologies are as advanced as the setup involving trapped ions interacting with lasers leibfried2003 ; haffner2008 ; bruzewicz2019 . Ion traps enable precise control over state initialization, dynamics, and system measurements. The possibility of arranging the ions in different spatial configurations, including 2- and 3d crystals, along with laser driving, opens the door to simulating a vast number of condensed matter systems blatt2012 ; lamata2014 ; joshi2020 ; monroe2021 . Additionally, the capability to engineer reservoirs renders the trapped ion systems ideal setups for investigating quantum thermodynamic cycles teixeira2022 . Understanding transport in such controllable circumstances can advance not only the theory of out-of-equilibrium systems and many-body physics, but also provide insights for the development of new technologies joshi1993 ; lepri2003 ; manzano2012 ; asadian2013 ; bermudez2013 ; freitas2015 ; nicacio2015 ; joulain2016 ; manzano2016 ; nicacio2016 ; guo2019 ; maier2019 ; meher2020 ; chen2022 ; yang2020 ; landi2022 ; yan2023 .
In this work, we are interested in exploring the transport response in varied coupling scenarios, linked to the variation of the laser intensity employed for manipulating the trapped ion. This variation of the coupling strength makes it imperative to use the dressed master equation (DME) formalism to achieve accurate, physical results scala2007 ; beaudoin2011 ; santos2014 ; levy2014 ; chiara2018 . In particular, we find an interesting relation approximately satisfied by the laser-ion coupling constant $\Omega$ and detuning $\delta$, as well as the trap frequency $\nu$. This relation reads $\Omega^{2}+\delta^{2}=(m\nu)^{2}$, and its fulfillment approximately gives the optimized current. Surprisingly, this relation is also related to the local maxima and minima of the leftover coherence in the steady state. Additionally, we show the controlled emergence of negative differential heat conductivity yomo2005 ; elste2006 ; li2006 ; benenti2009 ; he2010 in this system, a phenomenon characterized by a nonmonotonic behavior of the current with respect to temperature difference between the two reservoirs. Finally, we studied the asymmetric character of the current motz2018 ; simon2019 ; kalantar2021 with respect to the swap of reservoirs. In particular, we study how the rectification factor responds to controlled parameter changes.
The paper is organized as follows. In Sec. II we present the model we are going to consider, giving a brief review of the theoretical description of trapped ions, the treatment of the open quantum system via the dressed master equation (DME), and how to obtain heat current. We proceed to present some numerical results in Sec. III in a wide range of physical parameters. Finally, in Sec. IV we summarize the results and make some final remarks.
II Model
Figure 1: (Color online) Schematic illustration of the system in consideration. A trapped ion (gray, dashed-border rectangle) can be described by its motional (left red circle) and electronic (right blue circle) parts. These can be approximated, respectively, by a quantum harmonic oscillator of frequency $\nu$ and a two-state system of energy split $\delta$. The motional and electronic subsystems can be coupled with the use of a laser. This is represented by the coupling strength $\Omega$. By connecting thermal reservoirs at different temperatures $T_{M}$ and $T_{E}$ (red and blue, solid-border rectangles) to the motional and electronic parts, respectively, it is expected that the system will reach a stationary state with constant flux of energy given by the heat current $J_{SS}$ (purple arrow).
We are interested in studying the properties of a single trapped ion coupled to thermal reservoirs at different temperatures. A trapped ion can be effectively characterized by its internal electronic state and the position of its center of mass. Through the application of selection rules and appropriate detunings, the electronic subspace can be simplified into a two-level system. Similarly, by adjusting the electromagnetic trapping fields, the motion of the ion’s center of mass can be accurately portrayed as a harmonic oscillation around an equilibrium point along the trap axis leibfried2003 . The application of the laser induces coupling between the electronic and motional parts, and the Hamiltonian describing their interaction can be cast in the form moya-cessa2012
$H_{S}=\nu a^{\dagger}a+\frac{\delta}{2}\sigma_{z}+\frac{\Omega}{2}\bigl{[}\sigma_{+}\mathcal{D}(i\eta)+\sigma_{-}\mathcal{D}^{\dagger}(i\eta)\bigr{]},$ (1)
where $\nu$ is the ion’s vibrational frequency, $\delta:=\omega_{0}-\omega_{L}$ is the detuning between the electronic transition frequency $\omega_{0}$ and the laser frequency $\omega_{L}$, $\Omega$ is the coupling constant, and $\eta$ is the Lamb-Dicke (LD) parameter.
In this work, we explore the nonequilibrium consequences of introducing a temperature gradient across the ion under different coupling conditions. Typically, this temperature gradient is achieved through reservoir engineering. At a theoretical level, we capture the physics by subjecting both the electronic and motional degrees of freedom to independent Markovian baths at a specified temperature $T_{E}$ and $T_{M}$, respectively. Consequently, upon laser illumination, the two subsystems are expected to exchange energy through their laser-induced coupling, eventually settling into an asymptotic state. The energy current induced by the two thermal baths in this state is of particular interest to us. As we will see, the magnitude of the current exiting one reservoir equals the one entering the other in the asymptotic state. This scenario is depicted in Fig. 1.
The reservoirs are included in the dynamical description of the system by means of the bath Hamiltonian chen2022
$H_{B}=\sum_{\mu\in\{E,M\}}H_{B\mu}^{0}+V_{\mu},$ (2)
where
$H_{B\mu}^{0}=\sum_{k}\omega_{\mu k}b^{\dagger}_{\mu k}b_{\mu k}$ (3)
is the free Hamiltonian of the $\mu$-th reservoir, and
$V_{\mu}=A_{\mu}\otimes\sum_{k}g_{\mu k}(b_{\mu k}+b^{\dagger}_{\mu k})$ (4)
is the respective interaction term with the system $S$. We consider that the system-reservoir interaction is weak, and that reservoir correlations decay much faster than any significant time scale of the system. In these conditions, we can use the usual Born-Markov approximations, and it suffices to consider only up to linear interaction terms. Concerning the specific form of the operators in Eq. (4), we will be considering that the motional reservoir couples to the system with $A_{M}=a+a^{\dagger}$, while the electronic reservoir couples transversely, i.e., $A_{E}=\sigma_{x}$, which is a common choice in the context of energy exchange with the baths.
In the scope of the Born-Markov approximations, it is possible to show that the dynamics of the reduced system’s density operator, $\rho:=\mathrm{tr}_{B}\{\rho_{SB}\}$, is given by the dressed master equation (DME) breuer2002 ; scala2007 ; beaudoin2011
$\frac{d\rho}{dt}=-i[H_{S},\rho]+\sum_{\mu jk}\Gamma_{\mu jk}D_{jk}[\rho],$ (5)
where
$\Gamma_{\mu jk}=\begin{cases}\gamma_{\mu jk}\,\bar{n}_{\mu}(\omega_{jk})\,\lvert\langle k|A_{\mu}|l\rangle\rvert^{2},\text{ if }\omega_{jk}>0\\ \gamma_{\mu jk}\,[1+\bar{n}_{\mu}(-\omega_{jk})]\,\lvert\langle k|A_{\mu}|l\rangle\rvert^{2},\text{ if }\omega_{jk}<0\end{cases},$ (6)
and
$D_{jk}[\rho]=P_{k}\lvert j\rangle\langle j\rvert-\frac{1}{2}\lvert k\rangle\langle k\rvert\rho-\frac{1}{2}\rho\lvert k\rangle\langle k\rvert$ (7)
Here, $\lvert k\rangle$ is the eigenvector of $H_{S}$ such that $H_{S}\lvert k\rangle=E_{k}\lvert k\rangle$, $\omega_{jk}=E_{j}-E_{k}$, $P_{k}=\langle k|\rho|k\rangle$, $\bar{n}_{\mu}(\omega_{jk})=(e^{\omega_{jk}/T_{\mu}}-1)^{-1}$, and $\gamma_{\mu jk}$ are constants. These depend on particularities of the chosen bath. For simplicity, we will consider all of them to be equal, $\gamma_{\mu jk}=\gamma$.
Before moving forward, it is important to justify our use of the DME in place of the more common phenomenological, local master equation approach. In order to derive, e.g., a dissipator of the form
$D[\rho]=\gamma\biggl{(}a\rho a^{\dagger}-\frac{1}{2}\{a^{\dagger}a,\rho\}\biggr{)}$ (8)
from the system-bath interaction in Eq. (4), one has to neglect the coupling between the subsystems. In our case, between the electronic and motional parts. This is because the derivation of local master equations makes use of the free basis $\{\lvert e,n\rangle,\lvert g,n\rangle\}$. When the coupling becomes much stronger than the decay rates, the local master equation formalism starts to break. From a physical point of view, this can be seem as the system not being separable anymore. For example, the reservoir coupled with the motional part becomes indirectly coupled to the electronic part via the laser-induced interaction between them. From a mathematical point of view, this stems from the use of the dressed basis $\{\lvert k\rangle;H_{S}\lvert k\rangle=E_{k}\lvert k\rangle\}$ in the derivation of the DME in Eq. (5).
When placing a system between reservoirs at different temperatures, it is expected that the system does not reach an equilibrium state, but rather a time-independent, asymptotic stationary state. The total rate of energy change in this state is null, but it can be split into opposite-sign contributions which can be interpreted as a constant flux of energy leaving one reservoir and entering the other. This current is characterized by properties of both the baths and the system joshi1993 ; lepri2003 ; he2010 ; manzano2012 ; bermudez2013 ; freitas2015 ; nicacio2015 ; chen2022 ; landi2022 .
The starting point to find the heat current due to each of the reservoir is to notice that the total current reads
$J$ $\coloneqq-\frac{d}{dt}\langle H\rangle=-\mathrm{tr}\biggl{[}\frac{d\rho}{dt}H\biggr{]}$ (9)
$=-\sum_{\mu jk}\omega_{jk}\,\Gamma_{\mu jk}\,P_{k},$ (10)
where we have dropped the $S$ subscript from the system Hamiltonian in Eq. (1), and the minus sign indicates that the current is positive when energy leaves the system. For the steady state, which we denoted as $\rho_{SS}$, we have $d\rho_{SS}/dt=0$ in Eq. (5) and, consequently, $J=0$. From this, we can split the summation in $\mu\in\{E,M\}$ in Eq. (10) as $J=J_{E}+J_{M}=0$, from which we obtain the desired current for each reservoir in the steady state as
$J_{SS}\coloneqq J_{E}=\sum_{jk}\omega_{jk}\Gamma_{Ejk}P_{k}=-J_{M},$ (11)
where $J_{E}$ is the current between the electronic subsystem and its bath and $J_{M}$ is the analogous for the center of mass subsystem. As commented before, both currents are equal in magnitude in the steady state.
III Results
In this section we present some numerical simulations of the system’s response to the temperature gradient in several parameter regimes. All of the numerical results were obtained with the use of the QuTiP framework johansson2012 for Python. We started by calculating $\rho_{SS}$ from the DME in Eq. (5) and proceeded to calculate the current using Eq. (11). For all the results presented here, the Lamb-Dicke (LD) parameter was set to $\eta=0.05$ and the dimension of the bosonic mode was truncated at $N=30$. We experimented with other values of the LD parameter inside the LD regime, $\eta\ll 1$, and did not observe significant differences in the qualitative aspects of the results presented here. We also experimented with positive and negative values of the detuning $\delta$. However, the current magnitudes seem to only depend on $\lvert\delta\rvert$, with the current $J_{SS}$ varying less than $0.1\%$ between positive and negative detunings. Thus, without any loss of, we will present only the results for positive $\delta$.
Figure 2: (Color online) Steady-state magnitude of heat current $J_{SS}$ as a function of the detuning $\delta$ and the coupling strength $\Omega$. (a) $T_{E}=0.5\nu$ and $T_{M}=5\nu$. (b) $T_{E}=5\nu$ and $T_{M}=0.5\nu$.
In Fig. 2, the absolute value of the current is depicted as function of the detuning $\delta$ and coupling constant $\Omega$. In Fig. 2a, the electronic and motional reservoirs were set, respectively, at temperatures $T_{E}=0.5\nu$ and $T_{M}=5.0\nu$, while in Fig. 2b, we have the inverse scenario. Note that, in both cases, the maximum values of the current, the darker areas, form circular patterns in this $\Omega\times\delta$ space. Note also that the currents are asymmetric, having stronger maxima in the case where the electronic part is coupled to the hot reservoir. This asymmetry will be further explored shortly.
Figure 3: (Color online) Location of the local maxima of $\lvert J_{SS}\rvert$ (dashed purple lines) as a function of $\delta$ and $\Omega$. The circular sectors, $\delta^{2}+\Omega^{2}=(m\nu)^{2}$, $m=1,2,3$, are depicted as solid black lines. (a) $T_{E}=0.5\nu$ and $T_{M}=5\nu$. (b) $T_{E}=5\nu$ and $T_{M}=0.5\nu$.
In Fig. 3, it is possible to see that the local maxima of the heat current (dashed purple lines) tend to occur close to the circular sectors
$\delta^{2}+\Omega^{2}=(m\nu)^{2},$ (12)
with $m\in\mathbb{N}$, which are plotted as solid black lines. Interestingly enough, it was recently shown tassis2023 that, along the first circular sector, $\delta^{2}+\Omega^{2}=\nu^{2}$, the trapped ion Hamiltonian can be approximated, in a unitarily rotated frame, to that of the Jaynes-Cummings model. To the best of our knowledge, no analytical results or effective Hamiltonians are yet known for the other circular sectors, i.e., $m\neq 1$ in Eq. (12).
Another interesting result connected to Eq. (12) has to do with the residual coherence, i.e., coherence that still persists in the asymptotic state as the result of the interplay between coherent (Hamiltonian) and incoherent (thermal) dynamics. In order to quantify this quantum resource, we choose the relative entropy of coherence evaluated in the free basis =$\mathcal{F}=\{\lvert e,n\rangle,\lvert g,n\rangle\}$, which, in our case, reads baumgratz2014
$C(\rho)=S(\rho_{SS}^{\mathrm{diag}})-S(\rho_{SS}),$ (13)
where $S$ is the von Neumann entropy, and $\rho_{SS}^{\mathrm{diag}}$ has the same diagonal elements of $\rho_{SS}$ in the $\mathcal{F}$ basis and zeros in all other positions.
Figure 4: (Color online) Steady-state entropy of coherence $C$, calculated in the free Hamiltonian basis $\mathcal{F}$, as a function of $\delta$ and $\Omega$. (a) $T_{E}=0.5\nu$ and $T_{M}=5\nu$. (b) $T_{E}=5\nu$ and $T_{M}=0.5\nu$.
In Fig. 4, we present our results for the relative entropy of coherence, which, much like the current, displays circular patterns in the parameter space defined by $\nu$ and $\Omega$. It is intriguing that when the electronic component is coupled to the cold reservoir (Fig. 4a), the current peaks correspond to abrupt decreases in the relative entropy of coherence. Conversely, in the opposite scenario (Fig. 4b), current peaks are associated with peaks in residual coherence in the steady state. We emphasize that we are investigating the residual coherence left in the global system $S$ and not in the reduced states of the local electronic or motional degrees of freedom.
Figure 5: (Color online) Heat current $J_{SS}$ as a function of the temperature of the electronic reservoir, $T_{E}$. The temperature of the motional reservoir was $T_{M}=5\nu$, and the parameter values were $\delta=0.8\nu$ and $\Omega=1.5\nu$. Indicated by the dashed purple lines are the magnitude of the current maximum, and the current magnitude for the electronic reservoir at zero temperature, i.e., the largest temperature bias.
Having established sound connections between coherence and current, we now deepen our understanding of the current by studying how the trapped ion responds to varying the temperature gradient. It is usually expected that the magnitude of the current will be monotonic with the temperature bias, but this is not always the case. Some systems present what is called negative differential conductivity (NDC) yomo2005 ; elste2006 ; li2006 ; benenti2009 ; he2010 . In our system, we observe this phenomenon when we keep the temperature of the hot reservoir constant and we increase the temperature bias by lowering the temperature of the cold reservoir, as presented in Fig. 5. In this example, the motional reservoir was kept at temperature $T_{M}=5\nu$. Parameters used were $\delta=0.8\nu$ and $\Omega=1.5\nu$. Note the linear behaviour for temperatures larger than $T_{E}\approx 2\nu$, which is typical of the Fourier law joshi1993 ; lepri2003 ; manzano2012
$J\propto\Delta T.$ (14)
The linearity is broken for small temperatures of the reservoir coupled to the internal part, i.e., for stronger temperature bias. In particular, we observe nonmonotonicity for temperatures lower than $T_{E}\approx\nu$, which is the signature of NDC.
Figure 6: Ratio $Q$ as a function of $\delta$ and $\Omega$. For this plot, the motional and electronic reservoirs were set, respectively, at temperatures $T_{M}=5\nu$ and $T_{E}=0.5\nu$.
We can also investigate the appearance of this phenomenon for more general values of detuning $\delta$ and coupling constant $\Omega$. In order to do that, it is convenient to define a figure of merit for the NDC by means of the ratio
$Q:=\frac{\mathrm{max}\bigl{[}J_{SS}(T_{E})\bigr{]}}{J_{SS}(T_{E}=0)}.$ (15)
The graphic definitions of these quantities are depicted in Fig. 5. Basically, we have $Q=1$ when the current behaves monotonically with the temperature bias, and $Q>1$ otherwise. This is indeed the case displayed in Fig. 5. In Fig. 6 we present the values of the ratio $Q$ across the parameter space $\Omega\times\delta$. Again, here we set the motional reservoir at temperature $T_{M}=5\nu$, and calculated how $J_{SS}$ varies with $T_{E}$ for each pair $(\Omega,\delta)$, in order to compute $Q$. It can be seen that for small values of both $\delta$ and $\Omega$, we do not observe a significant amount of nonmonotonicity. For $\delta,\Omega\gtrsim 0.5\nu$ we observe an increase in the signatures of NDC. Again, the circular patterns tend to appear, but this time they are accompanied by oscillations in the value of $Q$. By fixing $\delta$, we observe oscillation in the value of $Q$ when approching and crossing the circular line. This is best noticed at the second circular sector.
The last aspect of transport we want to discuss in this setup is the onset of current rectification motz2018 ; simon2019 ; meher2020 ; kalantar2021 and its sensibility to variations in the system parameters. When the reservoirs are swapped, an alteration in the current direction is anticipated. Nevertheless, there might be an asymmetry in the current magnitudes, i.e., the current might flow more intensely in one direction. This is called current rectification. A quantifier of this asymmetry is found in the so-called rectification factor meher2020
$R\coloneqq\frac{\lvert J_{\rightarrow}\rvert-\lvert J_{\leftarrow}\rvert}{\max\{\lvert J_{\rightarrow}\rvert,\lvert J_{\leftarrow}\rvert\}},$ (16)
which is bounded by $\lvert R\rvert\leq 1$. When the currents flow symmetrically, we have $R=0$, while $\lvert R\rvert=1$ means complete rectification, i.e., the current flows only in one direction, like in a perfect diode.
Figure 7: (Color online) Rectification factor $R$ as a function of the detuning $\delta$ and the coupling strength $\Omega$. The hot and cold reservoirs were set, respectively, at temperatures $T_{H}=5\nu$ and $T_{C}=0.5\nu$
In Fig. 7, we present how the rectification factor depends on $\delta$ and $\Omega$. The temperature of the cold and hot reservoirs were set to $T_{C}=0.5\nu$ and $T_{H}=5\nu$, respectively. We define the current $J_{\rightarrow}$ as being the one when the cold reservoir is coupled to the electronic part, while $J_{\leftarrow}$ means the opposite, the electronic part being coupled to the hot reservoir. Once again, we observe the circular patterns show up. Despite being considerably asymmetric throughout the parameter ranges, we note that, close to the first circular sector, we have a clear bias towards negative rectification factors. In the second sector this also seem to be true, but with sharper changes in $R$ close to the circle line. In the third circle, however, the step size used in the numeric discretization did not give enough resolution for us to make a precise observation. In fact, as it can also be observed in the other figures, it appears that the variations in the features studied here become increasingly sharper for the outer circular sectors. This can ultimately constrain numerical studies for large integers $m$ in Eq. (12), since finer and finer discretization steps will be needed.
Overall, what we can observe in this last plot and the previous ones is that the system composed of a trapped ion interacting with a laser exhibits a remarkably rich behavior when subjected to a temperature gradient. The circular patterns and the entire phenomenology resulting from different coupling and detuning regimes make this system highly attractive for the investigation of the fundamentals of quantum mechanics and non-equilibrium thermodynamics neqt .
IV Final remarks
We investigated quantum heat transport through a trapped ion in several regimes of the coupling strength $\Omega$ and ion-laser detuning $\delta$. The heat current was driven by a reservoir coupled to the electronic part of the ion, and another coupled to the motional part. Since we considered couplings well into the strong regime, the use of the dressed master equation (DME) formalism became necessary. The current, then, was calculated for the steady state of the DME.
We found that the heat current forms an intriguing circular pattern in the $\delta\times\Omega$ space. The current maxima fall close to the circles $\delta^{2}+\Omega^{2}=(m\nu)^{2}$, for an integer $m$ multiple of the trap frequency $\nu$. Interestingly, it was recently shown tassis2023 that in the first circular sector, $\delta^{2}+\Omega^{2}=\nu^{2}$, it is possible to approximate, in a unitarily rotated frame, the trapped ion Hamiltonian to that of a Jaynes-Cummings model. This could assist future investigations of this system, since the JCM is an exactly solvable model. To the best of our knowledge, no such approximations are yet known for the other circular sectors.
Furthermore, we also observed that the residual coherence in the steady state, calculated with the canonical basis, also follows a similar circular pattern. When the hot reservoir is coupled to the motional part, we observed that current maxima correlate to sudden drops in the residual coherence. While, for the case of the hot reservoir coupled to the electronic reservoir, we had sudden peaks of leftover coherence when the current was maximal.
Other properties we investigated were differential heat conductivity and heat rectification. For the first, it was observed that, in the vast majority of parameter ranges considered, the system presents negative differential heat conductivity when one of the reservoirs was brought close to zero temperature. This means that, in those regions, the current presents nonmonotonic behaviour, decreasing as the temperature gradient increases. As for the heat rectification, a large asymmetry in the current was observed throughout the parameter space when the reservoirs were swapped. This may inspire applications such as thermal rectifiers and thermal diodes within this pivotal quantum technology setup. Finally, our study lays the foundation for further investigations into heat transport across controlled quantum systems.
Acknowledgements
T. T. acknowledges financial support from Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES, Finance Code 001). F. B. and F. L. S. acknowledge partial support from Brazilian National Institute of Science and Technology of Quantum Information (CNPq INCT-IQ 465469/2014-0). F. L. S. acknowledges partial support from Fundação de Amparo a Pesquisa do Estado de São Paulo (FAPESP) Process No. 2021/14135-1, CNPq (Grant No. 305723/2020-0), and CAPES/PrInt (Grant No. 88881.310346/2018-01).

### v2

@@
 Our focus lies in understanding how different laser coupling scenarios impact the system dynamics.
-As the laser intensity reaches a regime where the ion’s electronic and motional degrees of freedom strongly couple, traditional approaches using phenomenological models for thermal reservoirs become inadequate.
+As the laser intensity reaches a regime where the electronic and motional degrees of freedom of the ion couple strongly, traditional approaches using phenomenological models for thermal reservoirs become inadequate.
 Therefore, the adoption of the dressed master equation (DME) formalism becomes crucial, enabling a deeper understanding of how distinct laser intensities influence heat transport.
-Analyzing the heat current within the parameter space defined by detuning and coupling strength, we observe intriguing circular patterns which are influenced by the ion’s vibrational frequency and laser parameters, and reveal nuanced relationships between heat transport, residual coherence, and system characteristics.
-Our study also reveals phenomena such as negative differential heat conductivity and asymmetry in heat current flow, offering insights into the thermal properties of this essential quantum technology setup.
+Analyzing the heat current within the parameter space defined by detuning and coupling strength, we observe intriguing circular patterns which are influenced by the vibrational frequency of the ion and laser parameters, and reveal nuanced relationships between heat transport and coherence, as well as phenomena such as negative differential heat conductivity and heat rectification, offering insights into the thermal properties of this essential quantum technology setup.
 I Introduction
@@
 One of the efforts in this direction is that of understanding nonequilibrium processes, e.g., transport of energy, and how the response changes according to system or reservoir properties.
-On the other hand, few quantum technologies are as advanced as the setup involving trapped ions interacting with lasers leibfried2003 ; haffner2008 ; bruzewicz2019 .
+On the other hand, few quantum technologies are as advanced as the setup involving trapped ions interacting with lasers and cavity fields leibfried2003 ; haffner2008 ; bruzewicz2019 ; css .
 Ion traps enable precise control over state initialization, dynamics, and system measurements.
@@
 This relation reads $\Omega^{2}+\delta^{2}=(m\nu)^{2}$, and its fulfillment approximately gives the optimized current.
-Surprisingly, this relation is also related to the local maxima and minima of the leftover coherence in the steady state.
+Surprisingly, this relation is also related to local maxima or minima of the leftover coherence in the steady state.
 Additionally, we show the controlled emergence of negative differential heat conductivity yomo2005 ; elste2006 ; li2006 ; benenti2009 ; he2010 in this system, a phenomenon characterized by a nonmonotonic behavior of the current with respect to temperature difference between the two reservoirs.
 Finally, we studied the asymmetric character of the current motz2018 ; simon2019 ; kalantar2021 with respect to the swap of reservoirs.
-In particular, we study how the rectification factor responds to controlled parameter changes.
+In particular, we studied how the rectification factor responds to controlled parameter changes.
 The paper is organized as follows.
 In Sec.
-II we present the model we are going to consider, giving a brief review of the theoretical description of trapped ions, the treatment of the open quantum system via the dressed master equation (DME), and how to obtain heat current.
+II we present the model we are going to consider, giving a brief review of the theoretical description of trapped ions, the treatment of the open quantum system via the DME, and how to obtain the heat current.
 We proceed to present some numerical results in Sec.
@@
 II Model
-Figure 1: (Color online) Schematic illustration of the system in consideration.
-A trapped ion (gray, dashed-border rectangle) can be described by its motional (left red circle) and electronic (right blue circle) parts.
-These can be approximated, respectively, by a quantum harmonic oscillator of frequency $\nu$ and a two-state system of energy split $\delta$.
-The motional and electronic subsystems can be coupled with the use of a laser.
-This is represented by the coupling strength $\Omega$.
-By connecting thermal reservoirs at different temperatures $T_{M}$ and $T_{E}$ (red and blue, solid-border rectangles) to the motional and electronic parts, respectively, it is expected that the system will reach a stationary state with constant flux of energy given by the heat current $J_{SS}$ (purple arrow).
 We are interested in studying the properties of a single trapped ion coupled to thermal reservoirs at different temperatures.
 A trapped ion can be effectively characterized by its internal electronic state and the position of its center of mass.
-Through the application of selection rules and appropriate detunings, the electronic subspace can be simplified into a two-level system.
+Through the application of selection rules and appropriate detunings, the electronic subspace can be simplified into a two-level system leibfried2003 .
 Similarly, by adjusting the electromagnetic trapping fields, the motion of the ion’s center of mass can be accurately portrayed as a harmonic oscillation around an equilibrium point along the trap axis leibfried2003 .
 The application of the laser induces coupling between the electronic and motional parts, and the Hamiltonian describing their interaction can be cast in the form moya-cessa2012
-$H_{S}=\nu a^{\dagger}a+\frac{\delta}{2}\sigma_{z}+\frac{\Omega}{2}\bigl{[}\sigma_{+}\mathcal{D}(i\eta)+\sigma_{-}\mathcal{D}^{\dagger}(i\eta)\bigr{]},$ (1)
-where $\nu$ is the ion’s vibrational frequency, $\delta:=\omega_{0}-\omega_{L}$ is the detuning between the electronic transition frequency $\omega_{0}$ and the laser frequency $\omega_{L}$, $\Omega$ is the coupling constant, and $\eta$ is the Lamb-Dicke (LD) parameter.
+$H_{S}=\nu a^{\dagger}a+\frac{\delta}{2}\sigma_{z}+\frac{\Omega}{2}\bigl{[}\sigma_{+}e^{i\eta(a+a^{\dagger})}+\sigma_{-}e^{-i\eta(a+a^{\dagger})}\bigr{]},$ (1)
+where $\nu$ is the ion’s vibrational angular frequency, $\delta:=\omega_{0}-\omega_{L}$ is the detuning between the electronic transition frequency $\omega_{0}$ and the laser frequency $\omega_{L}$, $\Omega$ is the coupling constant, and $\eta$ is the Lamb-Dicke (LD) parameter.
+The operator $a$ ($a^{\dagger}$) annihilates (creates) a vibration quantum, $\sigma_{-}$ ($\sigma_{+}$) is the spin lowering (raising) operator, and $\sigma_{z}$ is the $z$ Pauli operator, which also acts on the spin subsystem.
+In this equation, and throughout the following, we adopt natural units where, in particular, the reduced Planck constant $\hbar$ (i.e., $h$ over $2\pi$) and the Boltzmann constant $k_{B}$ are set to one.
 In this work, we explore the nonequilibrium consequences of introducing a temperature gradient across the ion under different coupling conditions.
@@
 1.
+Figure 1: (Color online) Schematic illustration of the system in consideration.
+A trapped ion (gray, dashed-border rectangle) can be described by its motional (M) and electronic (E) subsystems, represented by the left red circle and the right blue circle, respectively.
+These can be approximated, respectively, by a quantum harmonic oscillator of frequency $\nu$ and a two-state system of energy split $\omega_{0}$.
+The motional and electronic subsystems can be coupled via a laser, with the coupling strength denoted by $\Omega$.
+By connecting thermal reservoirs at different temperatures $T_{M}$ and $T_{E}$ (red and blue, solid-border rectangles) to the motional and electronic parts, respectively, it is expected that the system will reach a stationary state with constant flux of energy given by the heat current $J_{SS}$ (purple arrow).
+From an implementation perspective, the thermal bath for the electronic degrees of freedom can be engineered by using dedicated lasers that promote the coupling of the target ion to other auxiliary ions and their common motion in the trap.
+As detailed in Duan , the spectral density of the resulting reservoir depends on various factors, including the ion number, the target ion location, the laser detuning relative to the motional sidebands, and the number of frequency components in the laser.
+Concerning the vibrational motion, laser cooling and heating can be used to implement effective thermal reservoirs.
+The interested reader can refer to teixeira2022 and the references therein.
+In Bermudez , this approach is considered to promote energy transport in phonon-mediated spin–spin interactions in crystals of trapped atomic ions.
+The physical picture here is that laser cooling couples the vibrational mode to an infinite number of photonic modes in the electromagnetic bath, thereby creating an effective thermal bath for the vibrational degrees of freedom.
 The reservoirs are included in the dynamical description of the system by means of the bath Hamiltonian chen2022
@@
 $V_{\mu}=A_{\mu}\otimes\sum_{k}g_{\mu k}(b_{\mu k}+b^{\dagger}_{\mu k})$ (4)
-is the respective interaction term with the system $S$.
+is the respective interaction term with the subsystem $\mu\in\{E,M\}$.
 We consider that the system-reservoir interaction is weak, and that reservoir correlations decay much faster than any significant time scale of the system.
-In these conditions, we can use the usual Born-Markov approximations, and it suffices to consider only up to linear interaction terms.
-Concerning the specific form of the operators in Eq.
-(4), we will be considering that the motional reservoir couples to the system with $A_{M}=a+a^{\dagger}$, while the electronic reservoir couples transversely, i.e., $A_{E}=\sigma_{x}$, which is a common choice in the context of energy exchange with the baths.
-In the scope of the Born-Markov approximations, it is possible to show that the dynamics of the reduced system’s density operator, $\rho:=\mathrm{tr}_{B}\{\rho_{SB}\}$, is given by the dressed master equation (DME) breuer2002 ; scala2007 ; beaudoin2011
+In these conditions, we can use the usual Born-Markov approximations.
+Regarding the specific form of the system operators in Eq.
+(4), we will be considering $A_{M}=a+a^{\dagger}$ and $A_{E}=\sigma_{x}$, which is a common choice to describe energy exchange with thermal baths.
+In this way, the dynamics of the system is given by the reduced density operator, $\rho:=\mathrm{tr}_{B}\{\rho_{SB}\}$, which obeys the dressed master equation (DME) breuer2002 ; scala2007 ; beaudoin2011
 $\frac{d\rho}{dt}=-i[H_{S},\rho]+\sum_{\mu jk}\Gamma_{\mu jk}D_{jk}[\rho],$ (5)
@@
 $D_{jk}[\rho]=P_{k}\lvert j\rangle\langle j\rvert-\frac{1}{2}\lvert k\rangle\langle k\rvert\rho-\frac{1}{2}\rho\lvert k\rangle\langle k\rvert$ (7)
-Here, $\lvert k\rangle$ is the eigenvector of $H_{S}$ such that $H_{S}\lvert k\rangle=E_{k}\lvert k\rangle$, $\omega_{jk}=E_{j}-E_{k}$, $P_{k}=\langle k|\rho|k\rangle$, $\bar{n}_{\mu}(\omega_{jk})=(e^{\omega_{jk}/T_{\mu}}-1)^{-1}$, and $\gamma_{\mu jk}$ are constants.
-These depend on particularities of the chosen bath.
-For simplicity, we will consider all of them to be equal, $\gamma_{\mu jk}=\gamma$.
-Before moving forward, it is important to justify our use of the DME in place of the more common phenomenological, local master equation approach. In order to derive, e.g., a dissipator of the form
+Here, $\lvert k\rangle$ is the eigenvector of $H_{S}$ such that $H_{S}\lvert k\rangle=E_{k}\lvert k\rangle$, $\omega_{jk}=E_{j}-E_{k}$, $P_{k}=\langle k|\rho|k\rangle$, $\bar{n}_{\mu}(\omega_{jk})=(e^{\omega_{jk}/T_{\mu}}-1)^{-1}$, and $\gamma_{\mu jk}$ are constants, which depend on particularities of the chosen baths.
+Unless explicitly stated, we consider equal relaxation rates $\gamma_{\mu jk}=\gamma$.
+Only when analyzing the phenomenon of current rectification do we consider a case with different such rates.
+Before proceeding, it is important to justify our use of the DME instead of the more common phenomenological, local master equation approach.
+To derive, for example, a local dissipator of the form scully
 $D[\rho]=\gamma\biggl{(}a\rho a^{\dagger}-\frac{1}{2}\{a^{\dagger}a,\rho\}\biggr{)}$ (8)
-from the system-bath interaction in Eq.
-(4), one has to neglect the coupling between the subsystems.
-In our case, between the electronic and motional parts.
-This is because the derivation of local master equations makes use of the free basis $\{\lvert e,n\rangle,\lvert g,n\rangle\}$.
-When the coupling becomes much stronger than the decay rates, the local master equation formalism starts to break.
-From a physical point of view, this can be seem as the system not being separable anymore.
-For example, the reservoir coupled with the motional part becomes indirectly coupled to the electronic part via the laser-induced interaction between them.
-From a mathematical point of view, this stems from the use of the dressed basis $\{\lvert k\rangle;H_{S}\lvert k\rangle=E_{k}\lvert k\rangle\}$ in the derivation of the DME in Eq.
-(5).
-When placing a system between reservoirs at different temperatures, it is expected that the system does not reach an equilibrium state, but rather a time-independent, asymptotic stationary state.
-The total rate of energy change in this state is null, but it can be split into opposite-sign contributions which can be interpreted as a constant flux of energy leaving one reservoir and entering the other.
-This current is characterized by properties of both the baths and the system joshi1993 ; lepri2003 ; he2010 ; manzano2012 ; bermudez2013 ; freitas2015 ; nicacio2015 ; chen2022 ; landi2022 .
-The starting point to find the heat current due to each of the reservoir is to notice that the total current reads
-$J$ $\coloneqq-\frac{d}{dt}\langle H\rangle=-\mathrm{tr}\biggl{[}\frac{d\rho}{dt}H\biggr{]}$ (9)
+for the vibrational motion, one must neglect its coupling to any other quantum system, except its own bath.
+In our case, this would correspond to neglecting the electronic degree of freedom.
+When the coupling constant $\Omega$ becomes much stronger than the decay rate $\gamma$, the local master equation formalism begins to break down, as the system is no longer separable.
+For instance, the reservoir coupled to the motional part becomes indirectly coupled to the electronic part via the laser-induced interaction between them.
+When a system is placed in contact with reservoirs at different temperatures, it is expected not to reach an equilibrium state but rather a time-independent, asymptotic stationary state.
+In this state, the total rate of energy change is zero, yet it can be split into contributions with opposite signs, interpreted as a constant energy flux from one reservoir to the other asadian2013 ; nicacio2016 .
+The starting point to find the heat current due to each of the reservoirs is to notice that the total current reads
+$J$ $\coloneqq-\frac{d}{dt}\langle H_{S}\rangle=-\mathrm{tr}\biggl{[}\frac{d\rho}{dt}H_{S}\biggr{]}$ (9)
 $=-\sum_{\mu jk}\omega_{jk}\,\Gamma_{\mu jk}\,P_{k},$ (10)
-where we have dropped the $S$ subscript from the system Hamiltonian in Eq.
-(1), and the minus sign indicates that the current is positive when energy leaves the system.
-For the steady state, which we denoted as $\rho_{SS}$, we have $d\rho_{SS}/dt=0$ in Eq.
-(5) and, consequently, $J=0$.
-From this, we can split the summation in $\mu\in\{E,M\}$ in Eq.
-(10) as $J=J_{E}+J_{M}=0$, from which we obtain the desired current for each reservoir in the steady state as
+where the minus sign indicates that the current is considered positive when energy leaves the system.
+We can now split the summation in $\mu\in\{E,M\}$ as $J=J_{E}+J_{M}$ to obtain the desired current for each reservoir.
+In particular, for the steady state denoted as $\rho_{SS}$, we have $d\rho_{SS}/dt=0$ and, consequently, $J=0$.
+In this scenario, we can write
 $J_{SS}\coloneqq J_{E}=\sum_{jk}\omega_{jk}\Gamma_{Ejk}P_{k}=-J_{M},$ (11)
-where $J_{E}$ is the current between the electronic subsystem and its bath and $J_{M}$ is the analogous for the center of mass subsystem.
-As commented before, both currents are equal in magnitude in the steady state.
+where $J_{E}$ is the current between the electronic subsystem and its bath and $J_{M}$ is the analogous current for the center-of-mass subsystem and its bath.
+As noted above, both currents are equal in magnitude in the steady state.
 III Results
-In this section we present some numerical simulations of the system’s response to the temperature gradient in several parameter regimes.
+This section presents numerical simulations of the system in Fig.
+1 responding to a temperature gradient across various parameter regimes.
 All of the numerical results were obtained with the use of the QuTiP framework johansson2012 for Python.
@@
 (11).
-For all the results presented here, the Lamb-Dicke (LD) parameter was set to $\eta=0.05$ and the dimension of the bosonic mode was truncated at $N=30$.
+For all the results presented here, the Lamb-Dicke (LD) parameter was set to $\eta=0.05$ and the dimension of the space of states of the motional bosonic mode (frequency $\nu$) was truncated at $N=30$ (truncated Fock basis).
 We experimented with other values of the LD parameter inside the LD regime, $\eta\ll 1$, and did not observe significant differences in the qualitative aspects of the results presented here.
@@
 However, the current magnitudes seem to only depend on $\lvert\delta\rvert$, with the current $J_{SS}$ varying less than $0.1\%$ between positive and negative detunings.
-Thus, without any loss of, we will present only the results for positive $\delta$.
-Figure 2: (Color online) Steady-state magnitude of heat current $J_{SS}$ as a function of the detuning $\delta$ and the coupling strength $\Omega$.
+Thus, without any generality loss, we will present only the results for positive $\delta$.
+Figure 2 shows simulation results for reservoirs at temperatures $T_{E}=0.5\nu$ and $T_{M}=5\nu$, with a fixed detuning of $\delta=0.8\nu$.
+The top panel displays the normalized current, $J_{SS}/(\gamma\nu)$, as a function of the coupling constant $\Omega$.
+Notably, the behavior of $J_{SS}$ is highly nonmonotonic with respect to $\Omega$, exhibiting a series of current peaks interspersed with regions of current suppression.
+The bottom panel shows the steady-state population $p_{n}$ in the truncated Fock basis used to represent the vibrational state in the simulations for a specific value of the Rabi frequency.
+It shows that the chosen number of elements in the basis, $N=30$, is sufficient to yield accurate numerical results.
+Figure 2: Top panel: Steady-state heat current $J_{SS}$ as a function of the coupling strength $\Omega$ for a fixed detuning $\delta=0.8\nu$.
+The reservoir temperatures were set to $T_{E}=0.5\nu$ and $T_{M}=5\nu$.
+Bottom panel: Steady-state population of the 30 states in the truncated Fock basis $|m\rangle$ (with $300\leq m\leq N=$ 30) for the bosonic motional mode, corresponding to the vibrational degree of freedom of the trapped ions (with frequency $\nu$).
+This plot was obtained with $\Omega=1.5\nu$, but similar results are obtained for all values of $\Omega$ considered in the top panel plot.
+In Fig.
+3, we extrapolate the previous simulation for several values of $\delta$ and present the normalized current as a function of both $\delta$ and $\Omega$.
+For Fig.
+3a, the electronic and motional reservoirs were set, respectively, at temperatures $T_{E}=0.5\nu$ and $T_{M}=5.0\nu$, while for Fig.
+3b, we have the inverse scenario.
+In both cases, the maximum values of the current, the darker areas, form circular patterns in the $(\Omega,\delta)$ space.
+Additionally, the currents are asymmetric, with stronger maxima when the electronic part is coupled to the hot reservoir.
+This asymmetry will be further explored shortly.
+Figure 3: (Color online) Magnitude of the steady-state heat current $J_{SS}$ as a function of the detuning $\delta$ and the coupling strength $\Omega$.
 (a) $T_{E}=0.5\nu$ and $T_{M}=5\nu$.
 (b) $T_{E}=5\nu$ and $T_{M}=0.5\nu$.
-In Fig.
-2, the absolute value of the current is depicted as function of the detuning $\delta$ and coupling constant $\Omega$.
-In Fig.
-2a, the electronic and motional reservoirs were set, respectively, at temperatures $T_{E}=0.5\nu$ and $T_{M}=5.0\nu$, while in Fig.
-2b, we have the inverse scenario.
-Note that, in both cases, the maximum values of the current, the darker areas, form circular patterns in this $\Omega\times\delta$ space.
-Note also that the currents are asymmetric, having stronger maxima in the case where the electronic part is coupled to the hot reservoir.
-This asymmetry will be further explored shortly.
-Figure 3: (Color online) Location of the local maxima of $\lvert J_{SS}\rvert$ (dashed purple lines) as a function of $\delta$ and $\Omega$.
+We now take a closer look at the appearance of the circular patterns.
+For each fixed value of $\delta$, we find the critical $\Omega$ that satisfies $dJ_{SS}/d\Omega=0$.
+The results are shown in Fig.
+4, where we can see that these local maxima tend to occur close to the circular sectors
+$\delta^{2}+\Omega^{2}=(m\nu)^{2},$ (12)
+where $m=1,2,\dots$ represents a natural number.
+This analysis confirms that, with the exception of a few isolated regions, particularly for $m=1$, such maxima of the current align closely with these circular sectors, as anticipated in Fig.
+3, confirming a significant relationship between the parameters.
+Interestingly, it was shown recently tassis2023 that along the first circular sector, $\delta^{2}+\Omega^{2}=\nu^{2}$, the trapped ion Hamiltonian can be approximated, in a unitarily rotated frame, to that of the Jaynes-Cummings model.
+While this result might appear to simplify the problem entirely, suggesting that the dynamics being fully described by the Jaynes-Cummings model would naturally lead to simple or intuitive analytical expressions for the heat current, this is not the case.
+Although the Jaynes-Cummings model is exactly solvable, its emergence in the present setup under the resonance $\delta^{2}+\Omega^{2}=\nu^{2}$ occurs within a nontrivial transformed frame tassis2023 .
+This picture relies on carefully chosen local and bipartite unitary transformations, leading to complex calculations even for simple observables.
+In particular, the expression for the heat current in Eq.
+(10) becomes a highly intricate form, comprising infinite sums of non-intuitive and cumbersome terms.
+Regarding the other circular sectors [Eq.
+(12) with $m\geq 2$], the underlying physical mechanism for their emergence is expected to be similar to the case of $m=1$, reported in tassis2023 .
+Specifically, these circular conditions $\delta^{2}+\Omega^{2}=(m\nu)^{2}$ are likely associated with the emergence of effective Hamiltonians that involve $m$ phonon transitions [$a^{m}$ and $(a^{\dagger})^{m}$], leading to generalizations of the Jaynes-Cummings model.
+This is consistent with the fact that $\nu$ denotes the vibration frequency (phonon mode).
+Deriving effective Hamiltonians for these cases requires identifying the appropriate unitarily transformed frames where the effective dynamics arise upon resonance, as discussed in tassis2023 for $m=1$.
+To the best of our knowledge, no such frames or effective Hamiltonians have yet been reported for $m\geq 2$, which makes this an interesting research problem by itself.
+[…]
+Figure 4: (Color online) Location of the local maxima (at constant $\delta$) of $\lvert J_{SS}\rvert$ (dashed purple lines) in the $(\Omega,\delta)$ space.
 The circular sectors, $\delta^{2}+\Omega^{2}=(m\nu)^{2}$, $m=1,2,3$, are depicted as solid black lines.
@@
 (b) $T_{E}=5\nu$ and $T_{M}=0.5\nu$.
-In Fig. 3, it is possible to see that the local maxima of the heat current (dashed purple lines) tend to occur close to the circular sectors
-$\delta^{2}+\Omega^{2}=(m\nu)^{2},$ (12)
-with $m\in\mathbb{N}$, which are plotted as solid black lines.
-Interestingly enough, it was recently shown tassis2023 that, along the first circular sector, $\delta^{2}+\Omega^{2}=\nu^{2}$, the trapped ion Hamiltonian can be approximated, in a unitarily rotated frame, to that of the Jaynes-Cummings model.
-To the best of our knowledge, no analytical results or effective Hamiltonians are yet known for the other circular sectors, i.e., $m\neq 1$ in Eq.
-(12).
-Another interesting result connected to Eq.
-(12) has to do with the residual coherence, i.e., coherence that still persists in the asymptotic state as the result of the interplay between coherent (Hamiltonian) and incoherent (thermal) dynamics.
-In order to quantify this quantum resource, we choose the relative entropy of coherence evaluated in the free basis =$\mathcal{F}=\{\lvert e,n\rangle,\lvert g,n\rangle\}$, which, in our case, reads baumgratz2014
-$C(\rho)=S(\rho_{SS}^{\mathrm{diag}})-S(\rho_{SS}),$ (13)
+Another interesting result connected to the circles in Eq.
+(12) relates to the coherence that still persists in the asymptotic state as the result of the interplay between coherent (Hamiltonian) and incoherent (thermal) dynamics.
+To quantify this quantum resource, we choose the relative entropy of coherence evaluated in the free basis =$\mathcal{F}=\{\lvert e,n\rangle,\lvert g,n\rangle\}$.
+It reads baumgratz2014
+$C(\rho_{SS})=S(\rho_{SS}^{\mathrm{diag}})-S(\rho_{SS}),$ (13)
 where $S$ is the von Neumann entropy, and $\rho_{SS}^{\mathrm{diag}}$ has the same diagonal elements of $\rho_{SS}$ in the $\mathcal{F}$ basis and zeros in all other positions.
-Figure 4: (Color online) Steady-state entropy of coherence $C$, calculated in the free Hamiltonian basis $\mathcal{F}$, as a function of $\delta$ and $\Omega$.
+In Fig.
+5, we present a comparison of the relative entropy of coherence $C(\rho_{SS})$ and the steady-state current $J_{SS}$ as a function of $\Omega$ and for fixed $\delta=0.8\nu$.
+In Fig.
+5a we have temperatures $T_{E}=0.5\nu$ and $T_{M}=5\nu$, while in Fig.
+5b, we have $T_{E}=5\nu$ and $T_{M}=0.5\nu$.
+We observe an intriguing correlation between current and coherence.
+In the first case, there appears to be a negative correlation, with coherence decreasing as current increases.
+In contrast, the second case exhibits a positive correlation, where peaks in the current align with peaks in coherence.
+Once again, we observe a pronounced asymmetry regarding the direction of the temperature gradient.
+Figure 5: (Color online) Magnitude of the steady-state current $J_{SS}$ and relative entropy of coherence $C$ of the steady state, calculated in the free Hamiltonian basis $\mathcal{F}$, as a function of $\Omega$ and for fixed $\delta=0.8\nu$.
 (a) $T_{E}=0.5\nu$ and $T_{M}=5\nu$.
 (b) $T_{E}=5\nu$ and $T_{M}=0.5\nu$.
+To further explore the relationship between current and coherence in the asymptotic state, we plot the coherence as a function of $\delta$ and $\Omega$, following the approach used for the current in Fig.
+3.
+The results are presented in Fig.
+6.
+Similar to the current, the figures reveal circular patterns in this parameter space.
+Notably, the correlation between current and coherence observed in Fig.
+5 appears to be consistent.
+Specifically, when the electronic component is coupled to the cold reservoir (Fig.
+6a), peaks in the current correspond to abrupt decreases in the relative entropy of coherence.
+Conversely, in the opposite scenario (Fig.
+6b), current peaks are associated with increases in the coherence present in the steady state.
+Figure 6: (Color online) Relative entropy of coherence $C$ of the steady state, calculated in the free Hamiltonian basis $\mathcal{F}$, as a function of $\delta$ and $\Omega$.
+(a) $T_{E}=0.5\nu$ and $T_{M}=5\nu$.
+(b) $T_{E}=5\nu$ and $T_{M}=0.5\nu$.
+Having established sound connections between coherence and current, we now deepen our understanding of the current by studying how the trapped ion behaves in response to varying temperature gradients.
+While it is generally expected that the current magnitude will increase monotonically with temperature bias, this is not always the case.
+Some systems exhibit what is known as negative differential conductivity (NDC) yomo2005 ; elste2006 ; li2006 ; benenti2009 ; he2010 .
+In the following, we set the temperature of the motional reservoir to $T_{M}=5\nu$ and vary the temperature of the reservoir coupled to the electronic part, $T_{E}$.
 In Fig.
-4, we present our results for the relative entropy of coherence, which, much like the current, displays circular patterns in the parameter space defined by $\nu$ and $\Omega$.
-It is intriguing that when the electronic component is coupled to the cold reservoir (Fig.
-4a), the current peaks correspond to abrupt decreases in the relative entropy of coherence.
-Conversely, in the opposite scenario (Fig.
-4b), current peaks are associated with peaks in residual coherence in the steady state.
-We emphasize that we are investigating the residual coherence left in the global system $S$ and not in the reduced states of the local electronic or motional degrees of freedom.
-Figure 5: (Color online) Heat current $J_{SS}$ as a function of the temperature of the electronic reservoir, $T_{E}$.
-The temperature of the motional reservoir was $T_{M}=5\nu$, and the parameter values were $\delta=0.8\nu$ and $\Omega=1.5\nu$.
-Indicated by the dashed purple lines are the magnitude of the current maximum, and the current magnitude for the electronic reservoir at zero temperature, i.e., the largest temperature bias.
-Having established sound connections between coherence and current, we now deepen our understanding of the current by studying how the trapped ion responds to varying the temperature gradient.
-It is usually expected that the magnitude of the current will be monotonic with the temperature bias, but this is not always the case.
-Some systems present what is called negative differential conductivity (NDC) yomo2005 ; elste2006 ; li2006 ; benenti2009 ; he2010 .
-In our system, we observe this phenomenon when we keep the temperature of the hot reservoir constant and we increase the temperature bias by lowering the temperature of the cold reservoir, as presented in Fig.
-5.
-In this example, the motional reservoir was kept at temperature $T_{M}=5\nu$.
-Parameters used were $\delta=0.8\nu$ and $\Omega=1.5\nu$.
-Note the linear behaviour for temperatures larger than $T_{E}\approx 2\nu$, which is typical of the Fourier law joshi1993 ; lepri2003 ; manzano2012
-$J\propto\Delta T.$ (14)
-The linearity is broken for small temperatures of the reservoir coupled to the internal part, i.e., for stronger temperature bias.
-In particular, we observe nonmonotonicity for temperatures lower than $T_{E}\approx\nu$, which is the signature of NDC.
-Figure 6: Ratio $Q$ as a function of $\delta$ and $\Omega$. For this plot, the motional and electronic reservoirs were set, respectively, at temperatures $T_{M}=5\nu$ and $T_{E}=0.5\nu$.
-We can also investigate the appearance of this phenomenon for more general values of detuning $\delta$ and coupling constant $\Omega$.
-In order to do that, it is convenient to define a figure of merit for the NDC by means of the ratio
-$Q:=\frac{\mathrm{max}\bigl{[}J_{SS}(T_{E})\bigr{]}}{J_{SS}(T_{E}=0)}.$ (15)
-The graphic definitions of these quantities are depicted in Fig.
-5.
-Basically, we have $Q=1$ when the current behaves monotonically with the temperature bias, and $Q>1$ otherwise.
-This is indeed the case displayed in Fig.
-5.
-In Fig.
-6 we present the values of the ratio $Q$ across the parameter space $\Omega\times\delta$.
-Again, here we set the motional reservoir at temperature $T_{M}=5\nu$, and calculated how $J_{SS}$ varies with $T_{E}$ for each pair $(\Omega,\delta)$, in order to compute $Q$.
-It can be seen that for small values of both $\delta$ and $\Omega$, we do not observe a significant amount of nonmonotonicity.
-For $\delta,\Omega\gtrsim 0.5\nu$ we observe an increase in the signatures of NDC.
-Again, the circular patterns tend to appear, but this time they are accompanied by oscillations in the value of $Q$.
-By fixing $\delta$, we observe oscillation in the value of $Q$ when approching and crossing the circular line.
-This is best noticed at the second circular sector.
-The last aspect of transport we want to discuss in this setup is the onset of current rectification motz2018 ; simon2019 ; meher2020 ; kalantar2021 and its sensibility to variations in the system parameters.
-When the reservoirs are swapped, an alteration in the current direction is anticipated.
-Nevertheless, there might be an asymmetry in the current magnitudes, i.e., the current might flow more intensely in one direction.
-This is called current rectification.
+7, we present results for a fixed detuning $\delta=0.8\nu$ and different values of the Rabi frequency $\Omega$.
+In all cases, we observe the onset of NDC, where the current no longer increases as the temperature gradient, $\Delta T\equiv T_{M}-T_{E}$, increases.
+However, for smaller temperature differences, the current exhibits the linear behavior $J\propto\Delta T$, consistent with Fourier’s Law.
+Similar NDC behavior is observed for different values of $\delta$ and $\Omega$, indicating a consistent response across these parameters.
+Additionally, when $\Delta T\approx 0$, the current behaves as expected, with heat flowing from the hot to the cold reservoirs, as indicated by the change in the signal of $J_{SS}$.
+Figure 7: (Color online) Heat current $J_{SS}$ as a function of the temperature bias $\Delta T\equiv T_{M}-T_{E}$.
+For each line, the coupling strength was set to: $\Omega=0.4\nu$ (dashed grey); $\Omega=0.8\nu$ (solid purple); $\Omega=1.5\nu$ (dotted black); and $\Omega=1.9\nu$ (dash-dotted cyan) The detuning was set to $\delta=0.8\nu$.
+The temperature of the motional reservoir was fixed at $T_{M}=5\nu$.
+Negative differential conductivity (NDC) is typically associated with the onset of nonlinearities.
+This well-established fact is exemplified by the well-known result that coupled harmonic oscillators cannot exhibit NDC li2006 ; he2010 .
+The results shown in Fig.(7) can thus be understood as arising from the inherent nonlinearity introduced by the laser-driven two-level system, which leads to the exponential function $\exp[\pm i\eta(a+a^{\dagger})]$ in the interaction part of the Hamiltonian, as given in Eq.(1).
+At sufficiently large temperature gradients, higher energy levels become populated, making the contribution of nonlinearity appreciable, leading to the observed NDC.
+The next and final aspect of transport we examine in this work is the onset of current rectification motz2018 ; simon2019 ; meher2020 ; kalantar2021 and its sensitivity to variations in the same parameters as before, which are tuned through the control of the laser — specifically, the detuning $\delta$ and the coupling constant $\Omega$.
+When the reservoirs are swapped, the current direction reverses, but an asymmetry in current magnitude may emerge, with a stronger flow in one direction.
+This can be noted by comparing the magnitudes of $J_{SS}$ in Figs.
+3a and 3b.
 A quantifier of this asymmetry is found in the so-called rectification factor meher2020
 $R\coloneqq\frac{\lvert J_{\rightarrow}\rvert-\lvert J_{\leftarrow}\rvert}{\max\{\lvert J_{\rightarrow}\rvert,\lvert J_{\leftarrow}\rvert\}},$ (14)
-which is bounded by $\lvert R\rvert\leq 1$.
-When the currents flow symmetrically, we have $R=0$, while $\lvert R\rvert=1$ means complete rectification, i.e., the current flows only in one direction, like in a perfect diode.
-Figure 7: (Color online) Rectification factor $R$ as a function of the detuning $\delta$ and the coupling strength $\Omega$.
-The hot and cold reservoirs were set, respectively, at temperatures $T_{H}=5\nu$ and $T_{C}=0.5\nu$
-In Fig.
-7, we present how the rectification factor depends on $\delta$ and $\Omega$.
-The temperature of the cold and hot reservoirs were set to $T_{C}=0.5\nu$ and $T_{H}=5\nu$, respectively.
-We define the current $J_{\rightarrow}$ as being the one when the cold reservoir is coupled to the electronic part, while $J_{\leftarrow}$ means the opposite, the electronic part being coupled to the hot reservoir.
-Once again, we observe the circular patterns show up.
-Despite being considerably asymmetric throughout the parameter ranges, we note that, close to the first circular sector, we have a clear bias towards negative rectification factors.
-In the second sector this also seem to be true, but with sharper changes in $R$ close to the circle line.
-In the third circle, however, the step size used in the numeric discretization did not give enough resolution for us to make a precise observation.
+where $J_{\rightarrow}$ is the current when the cold reservoir is coupled to the electronic part, and $J_{\leftarrow}$ is the current in the opposite scenario.
+From its definition in Eq.
+(14), we see that the rectification factor is bounded by $\lvert R\rvert\leq 1$.
+Specifically, $R=0$ indicates symmetric current flow, while $\lvert R\rvert=1$ represents complete rectification, where the current flows only in one direction, as in a perfect diode.
+Up to this point, we have considered equal relaxation rates for both reservoirs, $\gamma_{Ejk}=\gamma_{Mjk}=\gamma$, in Eq.(6).
+Next, we examine the effects on the rectification factor of introducing different such rates.
+In Fig.8, we present the rectification factor $R$ as a function of the coupling strength $\Omega$, for a fixed detuning $\delta=0.8\nu$.
+The three curves shown correspond to a fixed electronic reservoir relaxation rate $\gamma_{Ejk}=\gamma$, while varying the relaxation rates for the motional reservoir.
+Specifically, we considered the cases $\gamma_{Mjk}\equiv\gamma_{M}=0.5\gamma$ (dashed grey), $\gamma_{M}=\gamma$ (solid purple), and $\gamma_{M}=2\gamma$ (dash-dotted cyan).
+The locations of the circular sectors are marked by dotted vertical lines.
+It can be seen that the rectification factor is more affected by changes in the relaxation rates in some regions than in others.
+In particular, in the first circular sector (first dotted vertical line), an increase in the relaxation rate $\gamma_{M}$ leads to more negative values of $R$.
+Conversely, in the third circular sector, a decrease in $\gamma_{M}$ facilitates negative rectification.
+There are also small regions where the curves cross, which remain robust to small changes in the relaxation rates.
+This phenomenology suggests that the response of $R$ to variations in the thermal relaxation rates, combined with the flexibility offered by lasers of different intensities and detunings, opens up avenues for future investigations.
+These studies could focus on optimizing these parameters to achieve targeted values of the rectification factor.
+This approach could further broaden the potential applications of this system for thermal devices kalantar2021 and non-equilibrium thermodynamics neqt .
+Figure 8: (Color online) Rectification factor as a function of the coupling constant $\Omega$ for different decay rates.
+For each line, the decay rate of the reservoir coupled to the motional part was set to: $\gamma_{M}=0.5\gamma$ (dashed grey); $\gamma_{M}=\gamma$ (solid purple); and $\gamma_{M}=2\gamma$ (dash-dotted cyan).
+The decay rate for the electronic reservoir was always set to $\gamma_{E}=\gamma$ The hot and cold reservoirs were fixed, respectively, at temperatures $T_{H}=5\nu$ and $T_{C}=0.5\nu$.
+To gain a deeper understanding of how rectification depends on the parameters controlled by the laser properties, we once again fix the dissipation rates as $\gamma_{M}=\gamma_{E}=\gamma$ and present a simulation of the rectification factor for various values of $\delta$ and $\Omega$, as shown in Fig.
+9.
+Again, circular patterns tend to emerge across the parameter space, but now the phenomenology is more complex.
+We observe a tendency toward negative rectification factors near the first circular sector.
+This trend appears in the second sector as well, though with sharper variations in $R$ close to the circular boundary.
+When approaching the third circular sector, however, the step size used in the numeric discretization did not have enough resolution for us to make a precise observation.
 In fact, as it can also be observed in the other figures, it appears that the variations in the features studied here become increasingly sharper for the outer circular sectors.
@@
 (12), since finer and finer discretization steps will be needed.
-Overall, what we can observe in this last plot and the previous ones is that the system composed of a trapped ion interacting with a laser exhibits a remarkably rich behavior when subjected to a temperature gradient.
-The circular patterns and the entire phenomenology resulting from different coupling and detuning regimes make this system highly attractive for the investigation of the fundamentals of quantum mechanics and non-equilibrium thermodynamics neqt .
+Figure 9: (Color online) Rectification factor $R$ as a function of the detuning $\delta$ and the coupling strength $\Omega$.
+The hot and cold reservoirs were set, respectively, at temperatures $T_{H}=5\nu$ and $T_{C}=0.5\nu$
 IV Final remarks
-We investigated quantum heat transport through a trapped ion in several regimes of the coupling strength $\Omega$ and ion-laser detuning $\delta$.
-The heat current was driven by a reservoir coupled to the electronic part of the ion, and another coupled to the motional part.
-Since we considered couplings well into the strong regime, the use of the dressed master equation (DME) formalism became necessary.
-The current, then, was calculated for the steady state of the DME.
-We found that the heat current forms an intriguing circular pattern in the $\delta\times\Omega$ space.
-The current maxima fall close to the circles $\delta^{2}+\Omega^{2}=(m\nu)^{2}$, for an integer $m$ multiple of the trap frequency $\nu$.
-Interestingly, it was recently shown tassis2023 that in the first circular sector, $\delta^{2}+\Omega^{2}=\nu^{2}$, it is possible to approximate, in a unitarily rotated frame, the trapped ion Hamiltonian to that of a Jaynes-Cummings model.
-This could assist future investigations of this system, since the JCM is an exactly solvable model.
-To the best of our knowledge, no such approximations are yet known for the other circular sectors.
-Furthermore, we also observed that the residual coherence in the steady state, calculated with the canonical basis, also follows a similar circular pattern.
-When the hot reservoir is coupled to the motional part, we observed that current maxima correlate to sudden drops in the residual coherence.
-While, for the case of the hot reservoir coupled to the electronic reservoir, we had sudden peaks of leftover coherence when the current was maximal.
-Other properties we investigated were differential heat conductivity and heat rectification.
-For the first, it was observed that, in the vast majority of parameter ranges considered, the system presents negative differential heat conductivity when one of the reservoirs was brought close to zero temperature.
-This means that, in those regions, the current presents nonmonotonic behaviour, decreasing as the temperature gradient increases.
-As for the heat rectification, a large asymmetry in the current was observed throughout the parameter space when the reservoirs were swapped.
+We investigated quantum heat transport through a trapped ion in various regimes of the coupling strength $\Omega$ and ion-laser detuning $\delta$.
+The heat current was driven by one reservoir coupled to the electronic part of the ion and another coupled to the motional part.
+Given that we considered couplings well into the strong regime, the use of the dressed master equation (DME) formalism was necessary.
+The current was calculated for the steady state of the DME.
+We found that the heat current forms an intriguing circular pattern in the $(\delta,\Omega)$ space.
+The current maxima fall close to the circles $\delta^{2}+\Omega^{2}=(m\nu)^{2}$, for an integer $m$, where $\nu$ is the trap frequency.
+Furthermore, we observed that the coherence in the steady state, calculated with the decoupled basis, exhibits similar circular patterns, but with a notable distinction: it is not always the maximum that aligns with the circular lines.
+When the hot reservoir is coupled to the motional part, we observed that current maxima correlate with sudden drops in the residual coherence.
+Conversely, when the hot reservoir is coupled to the electronic part, we observed sudden peaks of leftover coherence when the current was maximal.
+Other properties we investigated included differential heat conductivity and heat rectification, where we observed a rich phenomenology as the coupling strength $\Omega$ was varied.
 This may inspire applications such as thermal rectifiers and thermal diodes within this pivotal quantum technology setup.
@@
 2021/14135-1, CNPq (Grant No.
-305723/2020-0), and CAPES/PrInt (Grant No.
+305723/2020-0) and (Grant No.
+313068/2023-2), and CAPES/PrInt (Grant No.
 88881.310346/2018-01).